# Particle detection performance and Geant4 simulation with low-cost CMOS technology

### v1

@@ 7 Conclusions
In this work, we implement a system for the data acquisition of the OmniVision OV5647 CMOS sensor with the help of a Raspberry Pi. We give precise information about the settings needed for the camera to take stable and sensitive frames for detecting particles. To reduce the experimental noise we subtract the average ADC value per pixel and apply a 5 sigma ADC threshold per pixel obtaining a reduction of 99.995% of active pixels.
We conclude from the $\chi^{2}_{\nu}$ of the analyzed distributions, for Sr90 and Cs137, that the simulation is in good agreement with the experimental data, considering the characteristics of the CMOS sensor. It is possible to carry out simulations of commercial sensors such as CMOS cameras if their physical specifications are precisely described. In addition, we can reproduce the correlation between variables (e.g. cluster size vs. maximum ADC per cluster) with a $\chi^{2}_{\nu}$ that shows good agreement for Sr90 and Cs137.
We analyzed how the detected source emission varies with the distance when the background noise is completely removed, finding that both the experimental data and the simulation follow an inverse square distribution, as expected.
When comparing the experimental data for the sources of Sr90 and Cs137 using their different parameters, we found that it is not possible with this experimental method to distinguish between the radioactive sources, nor using pure simulation.
A further step would be to identify the type of detected particles by the shape of the clusters and compare them between data and simulation.
Acknowledgements

### v2

@@ 7 Conclusions
In this work, we implement a system for the data acquisition of the OmniVision OV5647 CMOS sensor with the help of a Raspberry Pi. We give precise information about the settings needed for the camera to take stable and sensitive frames for detecting particles. We define a background reduction procedure, to eliminate experimental noise. We first subtract the average ADC value per pixel and apply a 5 sigma ADC threshold per pixel obtaining a reduction of 99.995% of active pixels. Then we remove clusters with a size of 5 pixels or fewer and clusters where one or more of its pixels is less than or equal to 100 ADC.
We conclude from the $\chi^{2}_{\nu}$ of the analyzed distributions, for Sr90 and Cs137, that the customized Geant4 simulation is in good agreement with the experimental data, considering the characteristics of the CMOS sensor. It is possible to carry out simulations of commercial sensors such as CMOS cameras if their physical specifications are precisely described. In addition, we can reproduce the correlation between variables (e.g. cluster size vs. maximum ADC per cluster) with a $\chi^{2}_{\nu}$ that shows good agreement for Sr90 and Cs137. Also, we find that both the experimental data and the simulation follow an inverse square distribution with the distance when the background noise is completely removed, as expected. Thus, this simulation, which has been cross-checked with data, can be used to test the feasibility of further particle detection ideas without the need to implement an experimental setup.
When comparing the experimental data for the Sr90 and Cs137 sources using the different analyzed parameters, we find that it is not possible to distinguish between the two radioactive sources with this experimental approach, nor through simulation alone. This limitation is related to the poor energy resolution of the sensor, which arises from the lack of a clear correlation between the deposited energy and the primary electron energy. However, given the accurate measurement of energy deposition corroborated by the simulation estimation, we have demonstrated that the sensor, once calibrated, is suitable for dosimetric measurements of source activities.
A refinement of this work  [^1] M. Bonnett Del Alamo et al. work in preparation. involves applying machine learning techniques to distinguish background clusters from signal ones. This approach could achieve comparable background rejection without sacrificing signal efficiency. Consequently, the ADC background cut could be loosened, which could enhance the sensor’s energy resolution, since in the low energy deposition region there is an anti-correlation between primary and deposited energy. Furthermore, the same method could be used for particle identification by analyzing cluster shapes and comparing them between data and simulation.
Acknowledgments